OpenStudy (anonymous):

0.25=0.9sin(pi x)+0.9 Can someone please show me why x=1.3

OpenStudy (whpalmer4):

Why do you think that \(x=1.3\)?

OpenStudy (whpalmer4):

I get \[x = -\frac{1}{\pi}\sin^{-1} (\frac{13}{18}) \approx -0.256879 \text{ radians}\]